Sol Summary – April 17th

Crew 245 Sol Summary Report 17-04-2022
SOL: 7
Summary Title: Martian Easter
Author’s name: Paolo Guardabasso

Mission status: Troubleshooting is ongoing for various experiments after some contingencies arose. The nice weather allows continuing EVA experimentation.

Sol Activity Summary: Calm day today: after an early morning EVA that was interrupted for low battery voltage in suits 3, the crew spent the day dedicating time to experiments but also relaxing.

At lunch, pasta with pesto was prepared for the crew. In the afternoon, another EVA was planned to collect biological samples and map an area with the drone. Luckily, the second lost sensor was found. In the meantime, some suit maintenance was performed, and some tests were run on the antenna assembly. Tonight, we’ll have a small celebration, sharing some more food from our countries’ traditions.

Look Ahead Plan: Tomorrow we’ll perform the first of two parts of the stress level experiment, where 2 crew members will have to explore and navigate an area with and without a digital map generated with the drone. Troubleshooting on the antenna will continue.

Anomalies in work: The morning EVA was interrupted because suit 3 was delivering low air. This was likely due to the fact that the charger connector detached from the suit and did not connect properly. The suit is now charging and we’ll test it later to make sure that the voltage is once again nominal. A proper test before EVA would’ve allowed us to spot the problem, so we prepared an EVA checklist, both for the night before and the 30 min before EVA.

Weather: Very warm day, with low winds. Temperature 7 (low) and 26 (high) C, Wind speed 19 (avg) and 35 (max) kph

Crew Physical Status: All nominal.

EVA: Apart from the mentioned suit 3 anomaly, EVAs went smoothly, and the crew explored a new area in the North.

Reports to be filled:
– Crew commander
– HSO report
– Operation report
– Mid-mission Research Report
– Journalist
– Picture of the day
– Some pictures
– EVA request (2)
– EVA report (2)

Support Requested: None at the moment.

Copyright © The Mars Society. All rights reserved. | Main Site